# Can AR Embedded Visualizations Foster Appropriate Reliance on AI in Spatial Decision-Making? A Comparative Study of AR X-Ray vs. 2D Minimap

### v3

@@ 1. Introduction
From high-stakes emergency evacuations (Tim2022interactiveEvacuation; Manfredi2024AMR) and crisis management (Launder2014ASI), to everyday situations like hurrying to a classroom or navigating an unfamiliar building under time pressure, people often have to rapidly choose among multiple spatial targets such as exits, areas to secure, and hallways. We refer to these decision-making tasks, which require integrating information about spatial environments with external data, as Spatial Decision-Making. Making a poor choice can waste time and resources, and in the worst cases, endanger lives (henig1996solving). Recent advancements in indoor sensing and artificial intelligence (AI) offer new avenues for supporting such decisions by visualizing real-time situational data and AI suggestions in a spatial context (XU2021Seethroughwalls; Nick2024OminiActions).
@@
However, the way this information is presented can significantly impact user behavior and decision quality (Gaba2024ModelUnFairPeopleCareOrNot). Existing methods (li2023autonomousgisnextgenerationaipowered; Xu2024EvaluatingLL) commonly employ 2D maps to display data and AI suggestions, requiring users to mentally map and connect these digital cues onto physical spaces. According to spatial cognition research (Klatzky1998AllocentricAE), this process, known as reference frame translation, can impose a substantial cognitive load. Under time pressure, users may struggle to critically assess AI outputs, resulting in inappropriate reliance (Swaroop2024AccuracyTime): either trusting flawed AI suggestions or rejecting correct ones. This concern is especially pressing in emergency contexts, where every second counts.
@@
Augmented Reality (AR) promises a more intuitive approach by directly visualizing data and AI suggestions onto the physical environment. These visualizations, often referred to as embedded visualizations (Willett2017EmbeddedDataRepresentations), can reduce cognitive load by eliminating the need for constant reference frame translation. Intuitively, embedded visualizations might foster more deliberate and appropriate reliance on AI. Yet, whether embedded visualizations genuinely mitigates inappropriate reliance on AI in spatial decision-making remains unclear, and little existing work has examined how embedded visualizations in AR affects human-AI reliance in spatial tasks. As AR and AI continue to mature and increasingly integrate (Hirzle2023WhenXRAndAIMeet), this work aims to provide researchers and practitioners with initial evidence on how embedded visualizations impact human-AI reliance in spatial decision-making.
@@
To this end, we present an empirical study (N=32) comparing an AR X-ray (Figure 1 Left) against a 2D Minimap (Figure 1 Right) in a time-pressured, AI-assisted spatial decision task. Following previous work on AR visual systems for indoor navigation (XU2021Seethroughwalls), we selected the X-ray view as our AR embedded visualization because it naturally embeds information within a large-scale indoor environment. In the tasks, participants needed to select one of four targets distributed across a two-floor building based on multiple criteria under time pressure. An AI with 75% accuracy was provided to suggest the optimal target, and both visualizations were experienced using an Apple Vision Pro (AVP)  [^1] https://www.apple.com/apple-vision-pro/, a state-of-the-art AR headset. This abstract task models time-pressured spatial choices with imperfect AI support, where over-reliance can hide errors and under-reliance can waste useful guidance.
@@ 2.1. Human-AI Decision-Making
Human-AI decision-making refers to collaborative decision processes where a human decision-maker interacts with an AI system to make a choice (2021BansalDoesTheWholeExceedItsParts). A common goal in these systems is to achieve complementary performance, where the human-AI team performs better than either the human or the AI alone (2021BansalDoesTheWholeExceedItsParts). However, empirical studies have repeatedly found that such complementary performance is rarely observed (beede2020human; carton2020feature). One widely reported barrier is inappropriate reliance — the inability of human users to correctly accept AI suggestions when they are right, or reject them when they are wrong (2021BansalDoesTheWholeExceedItsParts).
@@
Unlike trust in AI, which reflects users’ subjective feedback, reliance objectively measures whether a participant accepts or rejects the AI’s suggestion (2021BansalDoesTheWholeExceedItsParts; ma2023should), and has therefore been widely studied. To foster appropriate reliance, researchers have proposed providing explanations of the AI’s suggestions (schemmer2023AppRelianceByExplanation). Unfortunately, several studies found that explanations have often failed to deliver their intended benefits, but actually increase over-reliance – a phenomenon where users uncritically accept AI suggestions, even when they are suboptimal (2021BansalDoesTheWholeExceedItsParts). Drawn on the dual-process model of cognition (Chaiklin2012ThinkingFA; JSBTDualProcessTheory2013), initial efforts attributed this over-reliance to a lack of cognitive engagement with the explanation, because people often rely on heuristics (Type 1) to judge the AI’s suggestions, rather than engaging in deeper analysis (Type 2). Buçinca et al. (2021ToTrustOrToThink) introduced cognitive forcing functions that encourage users to think more critically about the AI’s suggestions. Vasconcelos et al. (vasconcelos2023explanationsreduceoverrelianceai) proposed a cost-benefit framework that formalizes when users choose to engage with AI explanations. Recently, Guo et al. (guo2024decision) developed a decision-theoretic framework to measure AI reliance by separating reliance behavior from cognitive limitations in signal interpretation.
@@
In the context of visualization research, automated methods to assist humans in making decisions with data have long been practiced. While a large body of work has focused on developing visual analytics systems (Saffo2024DesignSpaceOfIA), relatively fewer studies have investigated how visualizations affect human-AI collaboration. Among these studies, Yang et al. (yang2020visual) explored how visual explanations can foster appropriate trust in AI systems; Morrison et al. (morrison2023evaluating) investigated how different human explanation strategies might impact users’ reliance on AI in visual decision-making; Gaba et al. (Gaba2024ModelUnFairPeopleCareOrNot), Wang et al. (wang2022extending), and Wall et al. (wall2024trustjunk) each focused on how interface design influences users’ perceptions of model bias and performance, highlighting the powerful role of visual presentation in shaping trust in AI systems. Recently, Ha et al. (ha2024guided) found that in AI-guided visual analytics systems, users were more inclined to accept suggestions when completing a more difficult task despite the AI’s lower suggestion accuracy. Zhao et al. (zhao2023evaluating) and Reyes et al. (reyes2025trusting) both investigated how visualizing uncertainty in model outputs influences reliance and trust. Taken together, these studies underscore that visual presentation profoundly influences how users evaluate AI suggestions, rely on them, and ultimately make decisions.
@@ 2.2. Visualization-Aided Decision-Making
Within the field of visualization, research on decision-making typically falls into two major categories: understanding how humans use visualizations to make decisions (padilla2018decision) and design interactive systems (Oral_2023_ACriticalInquiry).
@@
The former body of work has drawn from theories in cognitive psychology to model and explain the mechanisms by which visualizations influence decision-making. For example, Padilla et al. (padilla2018decision) proposed a cognitive framework grounded in dual-process theories, illustrating how visualizations engage both intuitive (Type 1) and analytical (Type 2) reasoning processes in different contexts. Bancilhon et al. (bancilhon2023evaluating) expanded on this perspective by advocating for the integration of cognitive models in evaluating visualization effectiveness. A parallel line of work investigates how visualizations interact with human cognitive biases. Dimara et al. (dimara2018task) introduced a task-based taxonomy of cognitive biases relevant to visualization, while Wall et al. (wall2021left) explored how interaction traces might mitigate such biases. Bearfield et al. (bearfield2024same) further demonstrated that the same data visualization can elicit diverging interpretations. Research on uncertainty representations also plays an important role in understanding decision-making: Kale et al. (kale2020visual) showed that design choices, such as emphasizing means, can bias effect size judgments, while Fernandes et al. (fernandes2018uncertainty) found that quantile dotplots can significantly improve real-world decisions in uncertain scenarios.
@@
The latter body of work focuses on developing interactive visualization tools to support decision-making. Some of these systems are domain-agnostic and designed to facilitate specific decision processes, such as inspecting rankings (wall2017podium) and comparing alternatives based on multiple, often conflicting attributes (pajer2016weightlifter) in Multi-Criteria Decision-Making. Another line of work targets domain-specific decision contexts, such as medical treatment planning (kreiser2017decision) , urban policy (weng2018homefinder) , and more (chang2007wirevis). These systems typically incorporate domain-specific constraints to guide and contextualize the decision process. A recent survey by Oral et al. (Oral_2023_ACriticalInquiry) offers a comprehensive analysis of decision-focused visualization tools and highlights the opportunities and challenges in supporting all stages of the decision-making process.
@@
While decision-making is widely recognized as a core goal of visualization, Dimara and Stasko (Dimara2022WhereDoDecisionMakingHide) argue that it remains underrepresented in visualization research and task taxonomies. Brumar et al. (brumar2024typologydecisionmakingtasksvisualization) recently proposed a typology to better describe decision-making tasks in visualization. Despite this growing interest, decision support for spatial tasks remains an underexplored yet promising area (Goel2012DecisionmakingIA). Our study takes a first step toward investigating spatial decision support, aiming to provide a baseline and reference point for future research.
@@ 2.3. Spatial Cognition and Spatial Decision-Making
Spatial cognition encompasses the mental processes involved in perceiving, encoding, and reasoning about spatial environments (Tversky_2008_SpatialCognition). A fundamental aspect of spatial cognition is the use of reference frames—mental models that help individuals locate and orient themselves in space (Barbara2003StructureOfMentalSpace). Two commonly described reference frames are egocentric, where objects are located relative to the observer’s current position and orientation like “to my left or right”, and allocentric, where the environment is represented independently of the observer’s position, such as map-like or object-to-object relationships (Klatzky1998AllocentricAE). People switch flexibly between these frames, but doing so can introduce substantial cognitive load (Barbara2003StructureOfMentalSpace), particularly when tasks require precise mental transformations, such as rotating or scaling spatial representations in mind.
Spatial decision-making requires integrating information about physical environments with external data. To make such decisions with AI support, decision-makers must reconcile digital information with real-world contexts. With traditional 2D map-based visualizations, this process often involves switching between reference frames, which is cognitively demanding and can affect interpreting or verifying AI suggestions. By contrast, AR can embed information directly in the physical environment, thereby eliminating the spatial mapping overhead (Billinghurst2015ARSurvey). Thus, there is a sustained interest in leveraging AR systems to support spatial decision-making tasks.
@@ 2.4. AR, Situated, and Embedded Visualizations
Different from Virtual Reality (VR) (burdea2003virtual), AR can directly visualize information within physical environments, benefiting a broad range of real-world applications that require computational support (Kalkofen2009ComARVis). The visualization community has explored AR visualizations across multiple dimensions, including design patterns (lee2023design), techniques (Zhu_Tian_2024), systems (liu2025reality), empirical evaluations (quijano2024brushing), and applications (Lin_2021; lee2024sportify). We refer readers to the recent survey (Shin2024TheRealityofTheSituation) for more details.
More broadly, the concept of visualizing information in its physical context is known as Situated Visualization (Bressa2022SituatedVisualization), for which AR is one prominent realization among others like physicalization (Jansen2015DataPhysical). Depending on the spatial relationship between the visualization and its physical referent, Willett et al. (Willett2017EmbeddedDataRepresentations) distinguish between situated (co‑located in physical space) and embedded (overlaid or integrated with the referent). Based on this definition, our study compares two AR visualizations: a situated 2D Minimap and an embedded AR X-ray view. This comparison enables us to investigate whether embedding data directly into the physical referent can foster more appropriate reliance on AI by reducing the mental effort required to map data from a separate panel onto the physical environment. For clarity in the remainder of the paper, we refer to these two AR visualizations simply as Minimap and X-ray, respectively.
@@ 2.5. AR Visualizations for Spatial Decision-Making
Our particular interest lies in leveraging AR to support indoor spatial decision-making under time pressure, a crucial area with multiple applications such as emergency evacuation (Tim2022interactiveEvacuation; Sharma2020SAARIevacuation), first response operations (zhang2024OSTARFirstResponse), and security management (Su2024RASSAR). Despite its importance, research specifically addressing time-constraint spatial decision-making in AR remains limited. To the best of our knowledge, the most relevant work involves AR-based evacuation assistance, which provides users with dynamic, context-sensitive navigation cues to safely guide them toward exits during emergencies (zhang2024OSTARFirstResponse). For instance, Wächter et al. (Tim2022interactiveEvacuation) investigated AR-guided evacuation in buildings, showing that adapting AR visualizations to environmental hazards can improve evacuation efficiency and safety. Similarly, Sharma et al. (Sharma2020SAARIevacuation) developed AR modules to enhance situational awareness and cognitive mapping during evacuations in buildings.
Beyond emergency scenarios, a broadly related line of work explores AR-based indoor navigation, where users typically follow predefined routes rather than making spatial decisions. Most AR navigation techniques fall into two categories: annotation-based methods like arrows, and embedded visualizations. Annotations are arguably the most common AR navigation technique, as seen in applications like Google Maps [^2] https://maps.google.com. Numerous studies (Lee2022NavigationInsturctions; dong2021difference) have compared annotation-based AR navigation to traditional 2D Minimap across various contexts. Nonetheless, AR embedded visualizations offer benefits like a global view but have received less attention, partly due to technical challenges requiring the development of digital twins of real-world environments. Representative work in this direction includes Xu et al. (XUImroveIndoorWayfinding; XUARteambasedsearch) who conducted comparative studies of AR X-ray views (a type of AR embedded visualizations) for indoor wayfinding, demonstrating that egocentric perspectives significantly enhanced navigation efficiency, reduced cognitive load, and improved spatial awareness.

### v4

@@ 1. Introduction
From high-stakes emergency evacuations (Wächter et al., 2022; Manfredi et al., 2024) and crisis management (Launder and Perry, 2014), to everyday situations like hurrying to a classroom or navigating an unfamiliar building under time pressure, people often have to rapidly choose among multiple spatial targets such as exits, areas to secure, and hallways. We refer to these decision-making tasks, which require integrating information about spatial environments with external data, as Spatial Decision-Making. Making a poor choice can waste time and resources, and in the worst cases, endanger lives (Henig and Buchanan, 1996). Recent advancements in indoor sensing and artificial intelligence (AI) offer new avenues for supporting such decisions by visualizing real-time situational data and AI suggestions in a spatial context (Pu et al., 2021; Li et al., 2024).
@@
However, the way this information is presented can significantly impact user behavior and decision quality (Gaba et al., 2024). Existing methods (Li and Ning, 2023; Xu et al., 2024c) commonly employ 2D maps to display data and AI suggestions, requiring users to mentally map and connect these digital cues onto physical spaces. According to spatial cognition research (Klatzky, 1998), this process, known as reference frame translation, can impose a substantial cognitive load. Under time pressure, users may struggle to critically assess AI outputs, resulting in inappropriate reliance (Swaroop et al., 2024): either trusting flawed AI suggestions or rejecting correct ones. This concern is especially pressing in emergency contexts, where every second counts.
@@
Augmented Reality (AR) promises a more intuitive approach by directly visualizing data and AI suggestions onto the physical environment. These visualizations, often referred to as embedded visualizations (Willett et al., 2017; Zhu-Tian et al., 2023a), can reduce cognitive load by eliminating the need for constant reference frame translation. Intuitively, embedded visualizations might foster more deliberate and appropriate reliance on AI. Yet, whether embedded visualizations genuinely mitigates inappropriate reliance on AI in spatial decision-making remains unclear, and little existing work has examined how embedded visualizations in AR affects human-AI reliance in spatial tasks. As AR and AI continue to mature and increasingly integrate (Hirzle et al., 2023), this work aims to provide researchers and practitioners with initial evidence on how embedded visualizations impact human-AI reliance in spatial decision-making.
@@
To this end, we present an empirical study (N=32) comparing an AR X-ray (Figure 1 Left) against a 2D Minimap (Figure 1 Right) in a time-pressured, AI-assisted spatial decision task. Following previous work on AR visual systems for indoor navigation (Pu et al., 2021), we selected the X-ray view as our AR embedded visualization because it naturally embeds information within a large-scale indoor environment. In the tasks, participants needed to select one of four targets distributed across a two-floor building based on multiple criteria under time pressure. An AI with 75% accuracy was provided to suggest the optimal target, and both visualizations were experienced using an Apple Vision Pro (AVP)  [^1] https://www.apple.com/apple-vision-pro/, a state-of-the-art AR headset. This abstract task models time-pressured spatial choices with imperfect AI support, where over-reliance can hide errors and under-reliance can waste useful guidance.
@@ 2.1. Human-AI Decision-Making
Human-AI decision-making refers to collaborative decision processes where a human decision-maker interacts with an AI system to make a choice (Bansal et al., 2021). A common goal in these systems is to achieve complementary performance, where the human-AI team performs better than either the human or the AI alone (Bansal et al., 2021). However, empirical studies have repeatedly found that such complementary performance is rarely observed  (Beede et al., 2020; Carton et al., 2020). One widely reported barrier is inappropriate reliance — the inability of human users to correctly accept AI suggestions when they are right, or reject them when they are wrong (Bansal et al., 2021).
@@
Unlike trust in AI, which reflects users’ subjective feedback, reliance objectively measures whether a participant accepts or rejects the AI’s suggestion (Bansal et al., 2021; Ma et al., 2023), and has therefore been widely studied. To foster appropriate reliance, researchers have proposed providing explanations of the AI’s suggestions (Schemmer et al., 2023). Unfortunately, several studies found that explanations have often failed to deliver their intended benefits, but actually increase over-reliance – a phenomenon where users uncritically accept AI suggestions, even when they are suboptimal (Bansal et al., 2021). Drawn on the dual-process model of cognition (Chaiklin, 2012; Evans and Stanovich, 2013), initial efforts attributed this over-reliance to a lack of cognitive engagement with the explanation, because people often rely on heuristics (Type 1) to judge the AI’s suggestions, rather than engaging in deeper analysis (Type 2). Buçinca et al. (Buçinca et al., 2021) introduced cognitive forcing functions that encourage users to think more critically about the AI’s suggestions. Vasconcelos et al. (Vasconcelos et al., 2023) proposed a cost-benefit framework that formalizes when users choose to engage with AI explanations. Recently, Guo et al. (Guo et al., 2024) developed a theoretic framework to measure AI reliance by separating reliance behavior from cognitive limitations in signal interpretation.
@@
In the context of visualization research, automated methods to assist humans in making decisions with data have long been practiced. While a large body of work has focused on developing visual analytics systems (Saffo et al., 2024), relatively fewer studies have investigated how visualizations affect human-AI collaboration. Among these studies, Yang et al. (Yang et al., 2020) explored how visual explanations can foster appropriate trust in AI systems; Morrison et al. (Morrison et al., 2023) investigated how different human explanation strategies might impact users’ reliance on AI in visual decision-making; Gaba et al. (Gaba et al., 2024), Wang et al. (Wang et al., 2022), and Wall et al. (Wall et al., 2024) each focused on how interface design influences users’ perceptions of model bias and performance, highlighting the powerful role of visual presentation in shaping trust in AI systems. Recently, Ha et al. (Ha et al., 2024) found that in AI-guided visual analytics systems, users were more inclined to accept suggestions when completing a more difficult task despite the AI’s lower suggestion accuracy. Zhao et al. (Zhao et al., 2023) and Reyes et al. (Reyes et al., 2025) both investigated how visualizing uncertainty in model outputs influences reliance and trust. Taken together, these studies underscore that visual presentation profoundly influences how users evaluate AI suggestions, rely on them, and ultimately make decisions.
@@ 2.2. Visualization-Aided Decision-Making
In the field of visualization, research on decision-making typically falls into two categories: understanding how humans use visualizations to make decisions (Padilla et al., 2018a) and design interactive systems  (Oral et al., 2023).
@@
The former body of work has drawn from theories in cognitive psychology to model and explain the mechanisms by which visualizations influence decision-making. For example, Padilla et al. (Padilla et al., 2018a) proposed a cognitive framework grounded in dual-process theories, illustrating how visualizations engage both intuitive (Type 1) and analytical (Type 2) reasoning processes in different contexts. Bancilhon et al. (Bancilhon et al., 2023) expanded on this perspective by advocating for the integration of cognitive models in evaluating visualization effectiveness. A parallel line of work investigates how visualizations interact with human cognitive biases. Dimara et al. (Dimara et al., 2018) introduced a task-based taxonomy of cognitive biases relevant to visualization, while Wall et al. (Wall et al., 2021) explored how interaction traces might mitigate such biases. Bearfield et al. (Bearfield et al., 2024) further demonstrated that the same data visualization can elicit diverging interpretations. Research on uncertainty representations also plays an important role in understanding decision-making: Kale et al. (Kale et al., 2020) showed that design choices, such as emphasizing means, can bias effect size judgments, while Fernandes et al. (Fernandes et al., 2018) found that quantile dotplots can significantly improve real-world decisions in uncertain scenarios.
@@
The latter body of work focuses on developing interactive visualization tools to support decision-making. Some of these systems are domain-agnostic and designed to facilitate specific decision processes, such as inspecting rankings (Wall et al., 2017) and comparing alternatives based on multiple, often conflicting attributes (Pajer et al., 2016) in Multi-Criteria Decision-Making. Another line of work targets domain-specific decision contexts, such as medical treatment planning (Kreiser et al., 2017) , urban policy (Weng et al., 2018) , and more (Chang et al., 2007). These systems typically incorporate domain-specific constraints to guide and contextualize the decision process. A recent survey by Oral et al. (Oral et al., 2023) offers a comprehensive analysis of decision-focused visualization tools and highlights the opportunities and challenges in supporting all stages of the decision-making process.
@@
While decision-making is widely recognized as a core goal of visualization, Dimara and Stasko (Dimara and Stasko, 2022) argue that it remains underrepresented in visualization research and task taxonomies. Brumar et al. (Brumar et al., 2024) recently proposed a typology to better describe decision-making tasks in visualization. Despite this growing interest, decision support for spatial tasks remains an underexplored yet promising area (Goel et al., 2012). Our study takes a first step toward investigating spatial decision support, aiming to provide a baseline and reference point for future research.
@@ 2.3. Spatial Cognition and Spatial Decision-Making
Spatial cognition encompasses the mental processes involved in perceiving, encoding, and reasoning about spatial environments (Tversky, 2008). A fundamental aspect of spatial cognition is the use of reference frames—mental models that help individuals locate and orient themselves in space (Tversky, 2003). Two commonly described reference frames are egocentric, where objects are located relative to the observer’s current position and orientation like “to my left or right”, and allocentric, where the environment is represented independently of the observer’s position, such as map-like or object-to-object relationships (Klatzky, 1998). People switch flexibly between these frames, but doing so can introduce substantial cognitive load (Tversky, 2003), particularly when tasks require precise mental transformations, such as rotating or scaling spatial representations in mind.
Spatial decision-making requires integrating information about physical environments with external data. To make such decisions with AI support, decision-makers must reconcile digital information with real-world contexts. With traditional 2D map-based visualizations, this process often involves switching between reference frames, which is cognitively demanding and can affect interpreting or verifying AI suggestions. By contrast, AR can embed information directly in the physical environment, thereby eliminating the spatial mapping overhead (Billinghurst et al., 2015). Thus, there is a sustained interest in leveraging AR systems to support spatial decision-making tasks.
@@ 2.4. AR, Situated, and Embedded Visualizations
Different from Virtual Reality (VR) (Burdea and Coiffet, 2003), AR can directly visualize information within physical context, benefiting a broad range of applications that require computational support (Kalkofen et al., 2009). The visualization community has explored AR visualizations across multiple dimensions, including design patterns (Lee et al., 2023), techniques  (Zhu-Tian et al., 2024), interactions  (Liu et al., 2025; Zhu-Tian et al., 2023b), empirical evaluations (Quijano-Chavez et al., 2024; Pietschmann et al., 2023), and applications  (Lee et al., 2024; Lin et al., 2023). We refer readers to the recent survey (Shin et al., 2024) for more details.
More broadly, the concept of visualizing information in its physical context is known as Situated Visualization (Bressa et al., 2022), for which AR is one prominent realization among others like physicalization  (Jansen et al., 2015). Depending on the spatial relationship between the visualization and its physical referent, Willett et al. (Willett et al., 2017) distinguish between situated (co‑located in physical space) and embedded (overlaid or integrated with the referent). Based on this definition, our study compares two AR visualizations: a situated 2D Minimap and an embedded AR X-ray view. This comparison enables us to investigate whether embedding data directly into the physical referent can foster more appropriate reliance on AI by reducing the mental effort required to map data from a separate panel onto the physical environment. For clarity in the remainder of the paper, we refer to these two AR visualizations simply as Minimap and X-ray, respectively.
@@ 2.5. AR Visualizations for Spatial Decision-Making
Our particular interest lies in leveraging AR to support indoor spatial decision-making under time pressure, a crucial area with multiple applications such as emergency evacuation (Wächter et al., 2022; Sharma et al., 2020), first response operations (Zhang et al., 2024), and security management (Su et al., 2024). Despite its importance, research specifically addressing time-constraint spatial decision-making in AR remains limited. To the best of our knowledge, the most relevant work involves AR-based evacuation assistance, which provides users with dynamic, context-sensitive navigation cues to safely guide them toward exits during emergencies (Zhang et al., 2024). For instance, Wächter et al. (Wächter et al., 2022) investigated AR-guided evacuation in buildings, showing that adapting AR visualizations to environmental hazards can improve evacuation efficiency and safety. Similarly, Sharma et al. (Sharma et al., 2020) developed AR modules to enhance situational awareness and cognitive mapping during evacuations in buildings.
Beyond emergency scenarios, a broadly related line of work explores AR-based indoor navigation, where users typically follow predefined routes rather than making spatial decisions. Most AR navigation techniques fall into two categories: annotation-based methods like arrows, and embedded visualizations. Annotations are arguably the most common AR navigation technique, as seen in applications like Google Maps [^2] https://maps.google.com. Numerous studies (Lee et al., 2022; Dong et al., 2021) have compared annotation-based AR navigation to traditional 2D Minimap across various contexts. Nonetheless, AR embedded visualizations offer benefits like a global view but have received less attention, partly due to technical challenges requiring the development of digital twins of real-world environments. Representative work in this direction includes Xu et al. (Xu et al., 2024b, a) who conducted comparative studies of AR X-ray views (a type of AR embedded visualizations) for indoor wayfinding, demonstrating that egocentric perspectives significantly enhanced navigation efficiency, reduced cognitive load, and improved spatial awareness.